Calculate the arithmetic mean taking A=8.95 by (i) Short-cut Method and (ii) Coding Method for the following data.

Death Rate

3.5-4.4

4.5-5.4

5.5-6.4

6.5-7.4

7.5-8.4

8.5-9.4

9.5-10.4

10.5-11.4

11.5-12.4

12.5-13.4

13.5-14.4

F

1

4

5

13

12

19

13

10

6

4

1



1
Expert's answer
2021-05-07T14:31:37-0400

The continuous data distribution is as follows-





Here

"A=19, \sum d=-124, n=11"


(i) By short cut method-


Arithmetic Mean "=A+\dfrac{\sum d}{n}=19-\dfrac{124}{11}=19-11.27=7.73"


(ii) By coding method we will get the same mean as above. i.e "AM=7.73"